# AgriField3D: A Curated 3D Point Cloud and Procedural Model Dataset of Field-Grown Maize from a Diversity Panel

### v1

@@ 2.4 Data Processing
To separate individual plants, CloudCompare software was used. This process referred to as plant isolation, involved using the crop tool in CloudCompare to extract each plant from the circular arrangement. Identification and naming of the individual plants were facilitated by the paper labels placed in front of each plant during scanning. These labels ensured accurate association between physical plants and their digital representations. Once isolated, each plant was saved as an individual ‘.ply‘ file with proper naming conventions. This plant isolation process preserved the structural integrity of the plants while ensuring readiness for downstream analysis.
Following this, a quality control process was implemented. Extraneous elements such as walls, tables, and personnel inadvertently captured during scanning were removed. Each set of maize point clouds was manually inspected to verify its quality, ensuring that plant structures were intact, free from artifacts, and properly cropped. Multiple rounds of review were conducted to maintain accuracy and reliability across the dataset. Additionally, a metadata file was developed to document quality assessments and plant-specific characteristics, including the presence of tassels, maize ears, leaf count, and resolution. These metadata files provide a useful context for researchers, facilitating targeted analyses and ensuring the dataset’s suitability for diverse applications, including AI model training, phenotypic studies, and agricultural research.
@@ 3.1 Data Segmentation
The segmentation of maize plant point cloud data is a key contribution for enabling phenotypic analysis and machine learning applications. Automated graph-based segmentation methods were combined with manual checks in this procedure. We report precise separation of plant components, including leaves and stalks, while maintaining a high standard of quality.
@@ 3.1.1 Segmentation Workflow
First, the structural skeleton of the point cloud was extracted through a process known as skeletonization. This step utilized Laplacian-based contraction with adjusted parameters to generalize across different maize examples, effectively identifying the primary stalk and branching patterns. Following this, stalk node identification was performed to detect the top and bottom nodes of the maize stalk based on z-coordinate extremes, with a center bias parameter applied to prioritize median x and y coordinates.
@@ 3.1.3 Recoloring and Merging
Following segmentation, a recoloring and merging process was implemented. Each plant component was recolored to adhere to a consistent color scheme. For example, the first leaf in every plant was assigned the same color, followed by the second leaf, and so on, with the stalk uniformly colored across all samples. This uniformity enhanced the dataset’s usability and interoperability. Table 2 details the color ordering of the leaves and stalks.
@@
Table 2: Color Coding for Plant Components
@@
3.1.4 Segmented Dataset
The segmentation and refinement processes were applied to 520 plants.
Figure 4: Nine examples of segmented maize plant point clouds from the dataset, showcasing the color-coded segmentation data.
Figure 4 showcases nine examples of the segmented maize plant point clouds, highlighting the color-coded segmentation data among the 520 segmented plants in the dataset.
3.1.5 Procedural Model
Procedural modeling is a computational approach to generating structured 3D representations from raw data by applying algorithmic rules and mathematical functions. Unlike raw point clouds, which are unstructured and can be noisy or incomplete, procedural models create parametric surfaces that provide a compact, scalable, and editable representation of plant structures. This method enables precise reconstructions that preserve geometric accuracy while allowing for controlled variations, making it particularly useful for synthetic data augmentation, structural analysis, and AI-driven phenotyping applications. Procedural modeling is especially valuable for plant phenotyping, where capturing complex leaf curvature and growth patterns is essential for trait analysis and computational simulations. The segmented dataset of 520 maize plant point clouds was used as input for a procedural modeling approach developed in Hadadi et al. [34]. This method enabled the high-fidelity reconstruction of 3D surfaces that accurately capture the shape and structure of individual maize leaves. By leveraging this procedural model, each segmented plant was processed to NURBS-based surfaces. The procedural model ensured that the reconstructed surfaces were closely aligned with the original segmented point clouds. Figure 5 and Figure 6 illustrate examples of these reconstructions, showcasing the transition from raw point cloud data to the final NURBS surface representation.
Figure 5: Processing of B73 maize plant. (a) Raw data, (b) segmented data, and (c) final NURBS surface.
Figure 6: Processing of Mo17 maize plant. (a) Raw data, (b) segmented data, and (c) final NURBS surface.
These additional files complement the segmented point clouds, bridging raw data with high-fidelity surface models that can be utilized for diverse research applications.
Figure 7: NURBS surface reconstruction using the procedural model for four diverse maize plants.
Figure 7 illustrates the NURBS surface reconstructions derived from the procedural modeling approach applied to four distinct maize plants. The examples highlight the diversity in maize leaf structures across maize plant varieties. The process accurately captures the complex geometries of maize leaves, including their curvature and edges.
4 Data Downsampling
Several downsampling techniques were evaluated to identify the most suitable method for downsampling. We used Open3D[35] for point cloud processing including voxel grid downsampling, Poisson disk sampling, farthest point sampling, and random downsampling. In our implementation, we combine two downsampling steps. First, a voxel grid downsampling procedure is applied, which uniformly partitions the space and reduces the point count while preserving the overall structure of the plants. The algorithm iteratively increases the voxel size until the downsampled point cloud has just more than the target $N$ points. If the resulting point cloud still exceeds $N$, a fine-tuning step is performed using random sampling to exactly reach the desired count. This combined approach balances computational efficiency with data quality and ensures precise point counts. Additionally, to further enhance usability, we have provided a Python script to allow researchers to generate custom downsampled datasets with any desired point count $N$.
Figure 9: Comparison of histograms showing the mean point to point distances for the dataset downsampled to 100K points, 50K points and 10K points.
Figure 9 presents the histograms of mean nearest neighbor distances for the three downsampled versions. As the point count decreases, the mean point-to-point distances increase, reflecting the reduced resolution. These visualizations highlight the trade-off between dataset size and structural detail, enabling researchers to choose the resolution best suited to their computational and analytical needs.
4.1 Metadata
The metadata file contains the following columns:
  • IsConnected: Indicates whether the plant’s stem and leaves are fully connected with no gaps (Yes/No).
  • AreLeavesTouching: Specifies whether the leaves of the plant are touching each other (Yes/No).
  • NumberOfLeaves: Records the total number of leaves present on the plant.
  • HasEar: Indicates whether the plant has an ear (Yes/No).
  • HasTassel: Indicates whether the plant has a tassel (Yes/No).
  • Size: Specifies the number of points in the point cloud.
Column Name Value
TagName 0433
IsConnected Yes
IsEasyToCountLeaves Easy
AreLeavesTouching No
NumberOfLeaves 15
HasColor Yes
HasEar Yes
HasTassel Yes
Size 1,558,177 points
Comment This is a good plant.
The metadata not only provides a detailed description of each plant’s structural and phenotypic characteristics but also allows for targeted data selection. For example, researchers can quickly identify plants with a specific number of leaves or those with complete color information to suit their experimental requirements.
5 Potential Applications
Feature Extraction and Phenotypic Analysis:
Shape Representation and 3D Modeling:
Virtual Phenotyping and Computational Plant Modeling:
Agricultural Technology Development:
6 Conclusions
This study presents a high-resolution 3D point cloud dataset of maize plants. By capturing detailed structural features and providing comprehensive metadata, this dataset bridges a critical gap in agricultural research, enabling precise and scalable analyses of plant architecture, morphology, and genotype-phenotype relationships. The dataset’s segmentation, annotation, and standardized downsampling cater to a wide range of downstream applications, from feature extraction and shape representation to the development of machine learning models for plant phenotyping. Its versatility makes it a valuable tool not only for phenomics but also for agricultural technology development, such as precision farming and autonomous systems.
While the dataset represents a significant step forward in providing high-resolution 3D point clouds for plant phenomics, limitations must be acknowledged. A key limitation is the focus on a single growth stage. The dataset captures plants at vegetative maturity, providing valuable insights into their structural characteristics at this stage. However, it does not encompass the full developmental cycle of maize, such as seedling, flowering, or senescence stages. This restricts the scope of analyses to specific phenotypic traits relevant to vegetative maturity.
Future work will aim to address these limitations by expanding the dataset and incorporating additional metadata fields to enhance its utility. Specifically, future versions of the dataset will include data from multiple growth stages, enabling researchers to study dynamic phenotypic changes throughout the plant lifecycle. These additions will provide a more comprehensive resource for investigating developmental processes and genotype-phenotype interactions over time.
By providing this dataset openly to the research community, we aim to accelerate progress in plant science and crop improvement studies. The integration of detailed metadata further enhances its usability, allowing researchers to tailor their analyses to specific traits and research objectives. We encourage researchers to explore this dataset in diverse agricultural contexts, from studying plant responses to environmental conditions to improving crop yield and sustainability. Future expansions of this resource, including additional growth stages and environmental data, will further amplify its impact, fostering innovation and collaboration across the field.
Acknowledgments
This work is supported by the AI Research Institutes program [AI Institute for Resilient Agriculture (AIIRA), Award No. 2021-67021-35329] from the National Science Foundation and U.S. Department of Agriculture’s National Institute of Food and Agriculture. We acknowledge support from Iowa State University Plant Science Institute. YL was supported in part by NSF grants (DBI-1661475 and IOS-1842097) to PSS and others and a scholarship from the China Scholarship Council.

### v3

@@ 2.4 Data Processing
To separate individual plants, CloudCompare software was used. This process, referred to as plant isolation, involved using the crop tool in CloudCompare to extract each plant from the circular arrangement. Identification and naming of the individual plants were facilitated by the paper labels placed in front of each plant during scanning. These labels ensured accurate association between physical plants and their digital representations. Once isolated, each plant was saved as an individual ‘.ply‘ file with proper naming conventions. This plant isolation process preserved the structural integrity of the plants while ensuring readiness for downstream analysis.
Following this, a quality control process was implemented. Extraneous elements such as walls, tables, and personnel inadvertently captured during scanning were removed. Each set of maize point clouds was manually inspected to verify its quality, ensuring that plant structures were intact, free from artifacts, and properly cropped. Multiple rounds of review were conducted to maintain accuracy and reliability across the dataset. Additionally, a metadata file was developed to document quality assessments and plant-specific characteristics, including the presence of tassels, maize ears, and resolution. These metadata files provide a useful context for researchers, facilitating targeted analyses and ensuring the dataset’s suitability for diverse applications, including machine learning model training, phenotypic studies, and agricultural research.
@@ 3.1 Data Segmentation
The segmentation of maize plant point cloud data is a key contribution for enabling phenotypic analysis and ML applications. Automated graph-based segmentation methods were combined with manual checks in this procedure. We report precise separation of plant components, including leaves and stalks, while maintaining a high standard of quality.
@@ 3.1.1 Segmentation Workflow
First, the structural skeleton of the point cloud was extracted through a process known as skeletonization. This skeletonization was performed using the PC Skeletor library [39], which applies Laplacian-based contraction to produce a graph-based representation of the input point cloud. The parameters were adjusted to generalize across different maize examples, effectively identifying the primary stalk and branching patterns. Following this, stalk node identification was performed to detect the top and bottom nodes of the maize stalk based on z-coordinate extremes, with a center bias parameter applied to prioritize median x and y coordinates.
@@ 3.1.3 Recoloring and Merging
Following segmentation, a recoloring and merging process was implemented. Each plant component was recolored to adhere to a consistent color scheme. For example, the first leaf in every plant was assigned the same color, followed by the second leaf, and so on, with the stalk uniformly colored across all samples. This uniformity enhanced the dataset’s usability and interoperability. The color scheme includes labels for up to 16 leaves, which reflects the maximum number of segmented leaves in our segmented dataset. So we limit the color encoding to 16 leaf labels and the stalk. Table S1 in the Supplementary Materials details the color ordering of the leaves and stalks.
@@
3.1.4 Segmented Dataset
The segmentation and refinement processes were applied to 520 plants. Figure 2(b) showcases three examples of the segmented maize plant point clouds, highlighting the color-coded segmentation data among the 520 segmented plants in the dataset. Additional examples are provided in Figure S.2.
3.1.5 Procedural Model
Procedural modeling is a computational approach to generating structured 3D representations from raw data by applying algorithmic rules and mathematical functions. Unlike raw point clouds, which are unstructured and can be noisy or incomplete, procedural models create parametric surfaces that provide a compact, scalable, and editable representation of plant structures. This method enables precise reconstructions that preserve geometric accuracy while allowing for controlled variations, making it particularly useful for synthetic data augmentation, structural analysis, and phenotyping applications. Procedural modeling is especially valuable for plant phenotyping, where capturing complex leaf curvature and growth patterns is essential for trait analysis and computational simulations. The segmented dataset of 520 maize plant point clouds was used as input for a procedural modeling approach developed in Hadadi et al. [40]. This method enabled the high-fidelity reconstruction of 3D surfaces that accurately capture the shape and structure of individual maize leaves. By leveraging this procedural model, each segmented plant was processed to NURBS-based surfaces. The procedural model ensured that the reconstructed surfaces were closely aligned with the original segmented point clouds. Figure 4 illustrate examples of these reconstructions, showcasing the transition from raw point cloud data to the final NURBS surface representation.
Figure 4: Processing of two maize plants: B73 (top row) and Mo17 (bottom row). (a, d) Raw data, (b, e) segmented data, and (c, f) final NURBS surfaces.
These additional files complement the segmented point clouds, bridging raw data with high-fidelity surface models that can be utilized for diverse research applications. Figure 5 illustrates the NURBS surface reconstructions derived from the procedural modeling approach applied to four distinct maize plants. The examples highlight the diversity in maize leaf structures across maize plant varieties. The process accurately captures the complex geometries of maize leaves, including their curvature and edges.
Figure 5: NURBS surface reconstruction using the procedural model for five diverse maize plants.
4 Data Downsampling
Several downsampling techniques were evaluated to identify the most suitable method for downsampling. We used Open3D [41] for point cloud processing, including voxel grid downsampling (Voxel), Poisson disk sampling (Poisson), farthest point sampling (FPS), and random downsampling (Random). A comparison was performed for all four methods at three target resolutions—100k, 50k, and 10k points—to evaluate their runtime and output characteristics. As expected, Random downsampling was the fastest, although it lacked spatial consistency. FPS and Poisson Disk Sampling provided more uniform point distributions but incurred significantly higher computational costs. Voxel Grid sampling offered a balance, producing structured outputs with moderate runtime. Figure 7(a) qualitatively compares point distributions, supporting the effectiveness of our approach. Table S2 in Supplementary Materials summarizes the runtime performance of each method and
To balance efficiency and structural regularity, we adopted a two-step approach: Voxel Grid downsampling is first applied to preserve plant structure while reducing point count. The voxel size is iteratively increased until slightly over the target size $N$, followed by Random sampling to reach exactly $N$ points. This hybrid method ensures spatial uniformity with low computational cost.
Figure 7: (a) Qualitative comparison of downsampling strategies on representative point clouds. (b) Comparison of histograms showing the mean point to point distances for the dataset downsampled to 100K points, 50K points and 10K points.
We have provided a Python script to allow researchers to generate custom downsampled datasets with any desired point count $N$. Figure 7(b) presents the histograms of mean nearest neighbor distances for the three downsampled versions. As the point count decreases, the mean point-to-point distances increase, reflecting the reduced resolution. These visualizations highlight the trade-off between dataset size and structural detail, enabling researchers to choose the resolution best suited to their computational and analytical needs.
4.1 Metadata
The metadata file contains the following columns:
  • IsConnected: Indicates if the plant’s stem and leaves are connected with no gaps (Yes/No).
  • AreLeavesTouching: Specifies if the leaves of the plant are touching each other (Yes/No).
  • HasEar: Indicates whether the plant has an ear (Yes/No).
  • HasTassel: Indicates whether the plant has a tassel (Yes/No).
  • Size: Specifies the number of points in the point cloud.
Column Name Value
TagName 0433
IsConnected Yes
IsEasyToCountLeaves Easy
AreLeavesTouching No
HasColor Yes
HasEar Yes
HasTassel Yes
Size 1,558,177 points
Comment This is a good plant.
The metadata provides a detailed description of the structural and phenotypic characteristics of each plant (see Figure 8). In addition, it enables targeted data selection. For example, researchers can quickly identify plants with a specific number of leaves or those with complete color information to suit their experimental requirements.
5 Potential Applications
Feature Extraction and Phenotypic Analysis:
Shape Representation and 3D Modeling:
Virtual Phenotyping and Computational Plant Modeling:
Agricultural Technology Development:
5.1 Virtual Light Simulation with HELIOS
We performed a virtual light interception experiment using the HELIOS 3D modeling framework [37], to demonstrate a practical application of MaizeField3D. The goal was to evaluate the canopy-level light absorption across different maize varieties reconstructed from the dataset and validate the simulated results against field measurements.
Figure 9: Schematic of the HELIOS simulation. (a) A procedurally generated virtual field (left) based on MaizeField3D and the radiation simulation output (right) showing spatially varying absorbed PAR. (b) Hourly intercepted PAR (kW/m2) on the reconstructed field for a typical day.
Each 3D plant used in the simulation consisted of procedurally generated NURBS leaf surfaces derived from MaizeField3D, combined with stalk meshes reconstructed using the ball pivoting surface reconstruction algorithm in MeshLab. These reconstructions ensured realistic structural representations for simulation purposes. To generate genotype-specific virtual field plots, we created 42 unique structural variants for each of five selected genotypes. These variants were generated by applying randomized leaf rotations to the base plant, simulating natural morphological variability within a genotype. Specifically, for each variant, we rotated individual leaves around either the local stalk tangent (azimuthal rotation, up to ±10°) or a perpendicular axis (leaf tilt, up to ±5°). These augmented point clouds preserve genotype-level features while introducing realistic intra-genotype diversity [48]. We then assembled each virtual field plot using these 42 variants in a 3-row by 14-plant grid, following agronomic spacing standards (30 inches between rows, 6 inches between plants) [49]. This resulted in five simulated field plots, each corresponding to a different maize plant shown in Figure 5.
Figure 9(a) illustrates the visualization of the HELIOS simulation process. On the left, we show a procedurally generated virtual field constructed using MaizeField3D plant. On the right, we depict the HELIOS simulation output at one representative hour, where absorbed PAR is mapped across the canopy using color gradients. This highlights how structural details captured in MaizeField3D can be directly leveraged in physics-based light interaction models to analyze canopy-scale light distribution. We then used HELIOS to simulate diurnal PAR exposure over a full day (7:00 to 20:00). Figure 9(b) shows the intercepted PAR across the day for the same reconstructed field. The curve highlights the peak absorption around midday and the decrease toward morning and evening hours. The simulation was configured with fixed geographical coordinates (Ames, Iowa). Intercepted PAR was computed per hour using area-weighted radiation fluxes on plant surfaces. The corresponding total daily intercepted PAR values of the full field plots generated from the five plants (Figure 5) are reported in Table 2, revealing how structural variability influences light capture.
To validate the accuracy of these simulations, we replicated the experimental setup used for field validation by Zhou et al. [50]. As described in their work, field measurements were collected on August 7, 2020, 60 days after planting, during anthesis. The fraction of the PAR (fPAR) interception was estimated by subtracting ground-level measurements from above-canopy measurements and normalizing by the latter. In our simulation, we implemented an equivalent virtual sensing protocol by placing a horizontal sensor bar between the center rows of the simulated field. This allowed direct comparison of simulated fPAR values against field-measured values using the same spatial sampling strategy. We then compared the simulated fPAR values from the reconstructed fields against field-measured fPAR values for the same genotypes. As shown in Table 2, the simulated and measured fPAR values show strong agreement. This result supports the accuracy of our simulation pipeline in capturing canopy-scale light interception behavior across different genotypes.
Table 2: Total intercepted PAR energy and fractional PAR (fPAR) values for five simulated fields on a typical summer day. The plant names correspond to the labels shown in Figure 5.
Plant Total PAR (kWh/m2) Model fPAR Measured fPAR
a 42.44 0.96 0.96
b 31.97 0.87 0.95
c 28.26 0.86 0.76
d 30.84 0.64 0.65
e 30.26 0.84 0.94
6 Conclusions
This study presents a high-resolution 3D point cloud dataset of maize plants. By capturing detailed structural features and providing comprehensive metadata, this dataset bridges a critical gap in agricultural research, enabling precise and scalable analyses of plant architecture, morphology, and genotype-phenotype relationships. The dataset’s segmentation, annotation, and standardized downsampling cater to a wide range of downstream applications, from feature extraction and shape representation to the development of machine learning models for plant phenotyping.
While the dataset represents a step forward in providing high-resolution 3D point clouds for plant phenomics, limitations must be acknowledged. A key limitation is the focus on a single growth stage. The dataset captures plants at vegetative maturity, providing valuable insights into their structural characteristics at this stage. However, it does not encompass the full developmental cycle of maize, such as seedling, flowering, or senescence stages. This restricts the scope of analyses to specific phenotypic traits relevant to vegetative maturity. Future work will aim to address these limitations by expanding the dataset and incorporating additional metadata fields to enhance its utility. Specifically, future versions of the dataset will include data from multiple growth stages, enabling researchers to study dynamic phenotypic changes throughout the plant lifecycle. These additions will provide a more comprehensive resource for investigating developmental processes and genotype-phenotype interactions over time.
We also demonstrated a practical application of MaizeField3D by performing a virtual light interception experiment using the HELIOS 3D modeling framework. This use case illustrates how the dataset’s structurally accurate reconstructions can be leveraged in functional–structural plant modeling, enabling simulation of canopy-level light absorption under realistic field conditions. Such examples highlight the dataset’s utility for hypothesis testing and environmental interaction studies beyond static analysis.
By providing this dataset to the research community, we aim to accelerate progress in plant science and crop improvement studies. The integration of detailed metadata further enhances its usability, allowing researchers to tailor their analyses to specific traits and research objectives. We encourage researchers to explore this dataset in diverse agricultural contexts, from studying plant responses to environmental conditions to improving crop yield and sustainability. Future expansions of this resource, including additional growth stages and environmental data, will further amplify its impact, fostering innovation and collaboration across the field.
Data Availability
The MaizeField3D dataset is publicly available on the Hugging Face Datasets platform at https://huggingface.co/datasets/BGLab/MaizeField3D. It includes high-resolution point clouds, segmented plant models, metadata, and reconstructed outputs in STL and DAT formats. Full documentation and download instructions are provided on the project website: https://baskargroup.github.io/MaizeField3D/. The code for procedural NURBS surface generation used in this work is available at https://github.com/baskargroup/ProceduralMaize3D.
Acknowledgments
This work is supported by the AI Research Institutes program [AI Institute for Resilient Agriculture (AIIRA), Award No. 2021-67021-35329] from the National Science Foundation and U.S. Department of Agriculture’s National Institute of Food and Agriculture. We acknowledge support from the Iowa State University Plant Science Institute. YL was supported in part by NSF grants (DBI-1661475 and IOS-1842097) to PSS and others, and a scholarship from the China Scholarship Council.
Author CRediT Contributions
Conceptualization: AB, TZJ, AK, BG; Data curation: EK, MH, YL; Formal Analysis: MH, TZJ; Funding acquisition: AK, PSS, BG; Investigation: EK, MH, JG; Methodology: EK, MH, JG, AB, TZJ, AK, BG; Project administration: AK, PSS, BG; Resources: YL, PSS, BG; Software: EK, MH, JG, TZJ; Supervision: AB, TZJ, AK, PSS, BG; Validation: EK, MH, YL; Visualization: EK, MH; Writing – original draft: EK, MH, TZJ, AK, BG; Writing – review & editing: MH, TZJ, YL, AK, PSS, BG
Supplementary Materials
Figure S.1: Additional maize plant point clouds from the MaizeField3D dataset, showcasing the original data. These examples further illustrate the morphological diversity observed across maize plant varieties.
Figure S.2: Additional examples of segmented maize plant point clouds from the dataset.
Table S1: Color Coding for Plant Components
@@ Acknowledgments
Table S2: Runtime comparison (in seconds) of different downsampling strategies at three target sizes.
Method 100k 50k 10k
Random 0.03 ± 0.01 0.03 ± 0.01 0.02 ± 0.01
Voxel 0.95 ± 0.26 1.01 ± 0.32 1.11 ± 0.34
FPS 107.57 ± 38.24 55.99 ± 19.59 11.26 ± 3.93
Poisson 108.30 ± 37.22 57.01 ± 19.95 11.58 ± 4.09